if 5 men or 7 women earn 525 dollar per day.How much 7 would be 7 men or 13 women earn per day?




















Answer by Theo(13342)

525 / 5 = 105 per man per day.
525 / 7 = 75 per woman per day.

7 men would earn 7 * 105 = 735 per day.
13 women would earn 13 * 75 = 975 per day.